The Keypad and the viewing windowof the TI83 graphing calculator is rectangular. The length of the rectangle is 2cm more than twice the width, and the area of the rectangle is 144cm2. Find the length and the width.
Answer by sdmmadam@yahoo.com(530)

Let the length be L cms and
let the width be B cms
The length of the rectangle is 2cm more than twice the widt
That is L =2 B+2 ----(1)
the area of the rectangle is 144cm2
That is LB = 144 ----(2)
(2B+2)XB = 144
Dividing by 2
(B+1)B = 72
B^2+B-72 =0
(B+9)(B-8) = 0
(product = -72 and sum is 1 gives the quantiites to be 9 and (-8) )
(B+9) =0 gives B = -9 which is negative and hence not applicable as we are talking about the solids
(B-8) = 0 gives B = 8 ----(*)
Putting B = 8 in (2)
LB = 144
LX8 = 144
L=144/8 = 18
Answer: Length = 18 cms and Width = 8 cms
Verification: Length should be 2 greater than twice width
And ofcourse 18 is 2 greater than twice 8.
Hene our values are correct.